OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form. Thank you <3

OpenStudy (ghazi):

well lets say equation of line is y=mx+c now m=-1/4 because we have to take inverse of the slope of the given line cuz its perpendicular now y=-1/4 x+ c now to find c put x=2 and y=-3 -3=-1/4 * 2 + c -3=-1/2+c c=-3+1/2 c= -5/2 and your final equation is y=-1/4 x -5/2

OpenStudy (anonymous):

hmmm. ok! Thank you :)

OpenStudy (ghazi):

:D YW